Executive Cheese has issued debt with a market value of $110 million and has outstanding 16.00 million shares with a market price of $10 a share. It now announces that it intends to issue a further $52.00 million of debt and to use the proceeds to buy back common stock. Debtholders, seeing the extra risk, mark the value of the existing debt down to $62 million.

a. Calculate the market price of the stock following the announcement. (Round your answer to 2 decimal places.)

Price of the stock            $

b. How many shares can the company buy back with the $52.00 million of new debt that it issues? (Enter your answer in millions. Round your answer to 1 decimal place.)

Number of shares             million

c. What is the market value of the firm (equity plus debt) after the change in capital structure? (Enter your answer in millions. Round your answer to 2 decimal places.)

Market value            $  million

d. What is the debt ratio after the change in structure? (Round your answer to 2 decimal places.)

Debt ratio     
Answer #1

a. Price of the stock : $ 13

The market value of the firm's equity = $ 110 million - $ 62 million = $ 48 million.

Therefore the stock price = $ ( 160 + 48) / 16 = $ 13

b. Number of shares = 4 million.

The company can buy back = $ 52 million / $ 13 = 4 million shares.

c. Market value : $ 270 million

After the buy back of the shares, number of outstanding shares = 16 million - 4 million = 12 million.

Market value of equity = 12 million x $ 13 = $ 156 million.

Market value of debt = $ 62 million + $ 52 million = $ 114 million

Total market value = $ 156 million + $ 114 million = $ 270 million

d. Debt ratio after the change in capital structure : 0.42

Debt ratio after the change = 114 / 270 = 0.42